# AnyCap Project: A Unified Framework, Dataset, and Benchmark for Controllable Omni-modal Captioning

### v1

@@ 3.3 AnyCapEval
Figure 3: Evaluation methodology of AnyCapEval. (a) Examples demonstrating content evaluation via Keypoint Density (KPD) and style scoring rules. (b) KPD correlation analysis showing superior reliability. (c) Radar chart depicting performance gains from ACM integration.
We introduce AnyCapEval, a benchmark designed for systematic evaluation of controllability in omni-modal caption generation. Control requirements include two types. (1) Content assesses whether captions accurately follow control conditions and maintain semantic relevance. (2) Style evaluates consistency with human references in narrative manner, structure, and fluency.
Content evaluation. Given a reference caption $r$ and a user instruction $q$, we first annotate a key-point set $\mathcal{K}_{r,q}=\{k_{1},k_{2},\dots,k_{n}\}$ that exhaustively covers the information required by $q$. The candidate caption is denoted as $\hat{y}$. An automatic matcher based on GPT-4o [1] identifies the subset of key points present in $\hat{y}$ (Fig. 3(a) presents illustrative examples). To normalise for verbosity, we introduce the key-point density (KPD):
$\text{KPD}(\hat{y};{q})\;=\;\frac{N_{\text{match}}}{L_{\text{words}}(\hat{y})}\times 100,$ (3)
where $N_{\text{match}}$ denotes the count of key points from $\mathcal{K}_{r,q}$ present in $\hat{y}$, and $L_{\text{words}}(\hat{y})$is the total number of words in $\hat{y}$. KPD quantifies the effective information rate—matched key points per $100$ words.
While a naive alternative might simply count all extracted information points, KPD offers a more robust measure of effective information rate. Empirically, the Pearson correlation between this raw count ($R_{\text{all}}$) and human-judged relevance is weak, whereas the correlation for KPD in Eq. 3 is much stronger, as shown in Fig. 3(b). Length alone ($L_{\text{words}}$) already captures noise, so density provides a more stable, scale-invariant signal. Further details can be found in the Appendix B.
@@
Style evaluation. For style evaluation, GPT-4o compares the candidate caption $\hat{y}$ with the reference $r$ under the user instruction $q$ and assigns a discrete score $s(\hat{y})\in\{0,\dots,4\}$. The scoring criteria are carefully defined as follows: (0) severely deviates from $q$ or largely hallucinated/incorrect; (1) deviates from $q$ or contains many hallucinations; (2) slightly deviates from $q$ or contains some hallucination; (3) highly similar to $r$ and hallucination-free; (4) outperforms $r$ while remaining hallucination-free. (see Fig. 3(a) for examples).
This fine-grained scoring approach, guided by a detailed, structured prompt, supplies GPT-4o with the candidate $\hat{y}$, reference $r$, user instruction $q$, and explicit criteria along three axes—semantic similarity, hallucination severity, and stylistic conformity—for each level $0\to 4$. This tight rubric sharply reduces evaluation variance and subjective bias. Further details can be found in the Appendix B.
@@ 4 Experiments
4.1 Experimental Setup
We train the AnyCapModel (ACM) framework in two variants (2B and 8B) on our proposed AnyCapDataset (ACD), which comprises approximately 125K images, 100K videos, and 50K audio samples. Training is conducted using 32 NVIDIA A100 (80GB) GPUs, requiring just 6 hours for the 2B model and 21 hours for the 8B model. More training details are provided in the Appendix C.
4.2 Evaluation on Controllable Caption
Table 2: Results of controllable image captioning on AnyCapEval. Abbreviations like IPos. are different instruction types in Table 1. Content and style values are computed using KPD and style scores from Sec. 3.3, respectively. ACM significantly improves content accuracy and stylistic fidelity.
Model Content Style
@@ 4.2 Evaluation on Controllable Caption
 +ACM-2B 3.22 (+1.71) 4.49 (+1.06) 4.82 (+0.28) 4.00 (+1.32) 4.13 (+1.09) 2.27 (+0.14) 1.78 (-0.14) 2.38 (+0.44) 2.71 (+0.63) 2.67 (+0.15) 2.36 (+0.24)
 +ACM-8B 3.41 (+1.90) 4.82 (+1.39) 4.91 (+0.37) 3.89 (+1.21) 4.26 (+1.22) 2.33 (+0.20) 2.10 (+0.18) 2.56 (+0.62) 2.59 (+0.51) 2.83 (+0.31) 2.46 (+0.34)
Qwen2.5-VL-7B 1.51 3.52 5.18 3.03 3.31 2.00 2.02 2.50 2.14 2.63 2.22
 +ACM-2B 3.38 (+1.87) 4.42 (+0.90) 4.94 (-0.24) 4.68 (+1.65) 4.36 (+1.05) 2.40 (+0.40) 2.04 (+0.02) 2.50 (+0.00) 2.94 (+0.80) 2.72 (+0.09) 2.50 (+0.28)
 +ACM-8B 3.47 (+1.96) 4.45 (+0.93) 5.84 (+0.66) 4.34 (+1.32) 4.53 (+1.22) 2.36 (+0.36) 2.20 (+0.18) 2.82 (+0.32) 2.88 (+0.74) 2.74 (+0.11) 2.56 (+0.34)
Table 3: Main results of controllable video captioning on AnyCapEval. (See Appendix C for details) ACM substantially enhances content adherence and style fidelity, particularly in complex video tasks.
Model Content Style
@@
 +ACM-2B 4.67 (+1.59) 6.32 (+1.99) 3.19 (-0.25) 6.49 (+3.33) 4.50 (+1.95) 5.07 (+1.55) 1.84 (+0.45) 1.80 (+0.03) 2.48 (+0.57) 2.39 (+0.05) 2.19 (+0.26)
 +ACM-8B 4.95 (+1.87) 6.39 (+2.06) 7.03 (+3.59) 5.66 (+2.50) 4.95 (+2.40) 5.73 (+2.21) 2.00 (+0.61) 1.88 (+0.11) 2.20 (+0.29) 2.61 (+0.27) 2.24 (+0.31)
Qwen2.5-VL-7B 2.88 4.20 3.29 2.24 2.93 3.54 1.63 1.65 2.40 2.29 2.10
 +ACM-2B 3.39 (+0.51) 5.97 (+1.77) 6.23 (+2.94) 6.30 (+4.06) 4.48 (+1.55) 5.25 (+1.71) 1.66 (+0.03) 1.82 (+0.17) 2.30 (-0.10) 2.55 (+0.26) 2.16 (+0.06)
 +ACM-8B 4.03 (+1.15) 6.15 (+1.95) 6.58 (+3.29) 6.04 (+3.80) 4.82 (+1.89) 5.55 (+2.01) 1.92 (+0.29) 1.98 (+0.33) 2.45 (+0.05) 2.55 (+0.26) 2.31 (+0.21)
Table 4: Results on audio captioning using AnyCapEval. The table shows that ACM consistently boosts content and style metrics, demonstrating its strong effectiveness and adaptability even in low-resource audio modality environments.
Model Content Style
@@
 +ACM-2B 1.55 (+0.13) 1.37 (+0.35) 1.32 (+0.64) 0.92 (+0.49) 1.21 (+0.50)
 +ACM-8B 1.48 (+0.06) 1.23 (+0.21) 1.29 (+0.61) 1.13 (+0.70) 1.22 (+0.51)
Table 5: Comparison across Training Methods.
Base Model SFT DPO SC ACM
InternVL2.5-2B -2.8% +7.0% +2.8% +16.9%
Table 6: Training Data Ratios Comparison.
Base Model 1:1:1 1:2:2 2:1:2 2:2:1
InternVL2.5-8B +17.9% +17.2% +16.5% +21.1%
Figure 4: Performance across Data Ratios.
@@
We evaluate the impact of integrating ACM with various pretrained base models (proprietary models, e.g., GPT-4o [1]; open-source models, e.g., InternVL [32], Qwen2.5-VL [62], MiniCPM-o [63]) using our AnyCapEval benchmark across the image, video and audio. As shown in Tables 2, 3 and 4.2, adding ACM leads to robust improvements in controllability across all modalities and base models, particularly enhancing content fidelity. Fig. 3(c) visualizes these gains for GPT-4o and InternVL2.5-8B. ACM also demonstrates notable strength in improving control for video captioning tasks.
@@
Model-centric analysis. The results consistently demonstrate the effectiveness of ACM. Both the 2B and 8B variants significantly improve controllability scores over the unassisted base models across all modalities. The larger ACM-8B generally yields greater improvements. Furthermore, integrating ACM substantially elevates the performance of open-source models. With the ACM-8B, several open-source models achieve controllability scores on AnyCapEval that are comparable or superior to strong unassisted proprietary baselines (e.g. ACM-8B elevates the performance of InternVL2.5, an open-source model with initially modest results in the image style category, from $\sim 2.1$ to $\sim 2.5$. This enhanced score notably surpasses the $\sim 2.3$ achieved by GPT-4o in the same category).
@@
Task-centric analysis. Content-related improvements consistently exceed those on style, likely because content controls offer more explicit learning signals, whereas stylistic constraints are inherently subjective and harder to enforce. Visual tasks outperform audio counterparts, potentially due to the richer supervision and higher information density in vision-language pretraining. ACM exhibits variable impact across settings: in high-resource dimensions (e.g., image appearance), it refines already strong captions; in low-resource, complex scenarios (e.g., video actions or audio events), it must synthesize missing content, yielding larger relative gains. Rare minor regressions on specific metrics are attributed to inherent optimization trade-offs in balancing multifaceted controllability.
@@ 4.3 Evaluation on Public Benchmarks
Table 7: Performance on MIA-Bench. Abbreviation: desc.(description), ment.(mention), len.(length limit), persp.(perspective), gen.(genre), avg.(average).
Model desc. ment. len. persp. gen. avg.
GPT-4o 90.3 87.9 90.5 85.2 91.6 89.1
 +ACM-2B 89.1 (-1.2) 88.9 (+1.0) 90.3 (-0.2) 90.0 (+4.8) 92.2 (+0.6) 90.1 (+1.0)
 +ACM-8B 90.4 (+0.1) 89.4 (+1.5) 89.9 (-0.6) 89.2 (+4.0) 92.9 (+1.3) 90.3 (+1.2)
InternVL2.5-8B 85.2 79.0 82.6 72.5 85.2 80.9
 +ACM-2B 85.7 (+0.5) 79.6 (+0.6) 85.2 (+2.6) 77.5 (+5.0) 85.7 (+0.5) 82.7 (+1.8)
 +ACM-8B 86.9 (+1.7) 79.0 (+0.0) 86.1 (+3.5) 88.9 (+16.4) 86.3 (+1.1) 85.4 (+4.5)
Yi-VL-34B 71.7 52.0 55.8 60.0 59.8 59.9
 +ACM-2B 71.8 (+0.1) 54.2 (+2.2) 61.9 (+6.1) 61.7 (+1.7) 61.0 (+1.2) 62.1 (+2.2)
 +ACM-8B 79.5 (+7.8) 66.5 (+14.5) 67.1 (+11.3) 63.0 (+3.0) 75.8 (+16.0) 70.4 (+10.5)
Table 8: Performance on VidCapBench. Abbreviation: acc.(accuracy), pre.(precision), cov.(coverage), con.(conciseness).
Model acc. pre. cov. con.
GPT-4o 15.6 58.9 87.8 6.3
 +ACM-2B 15.4 (-0.2) 59.1 (+0.2) 87.1 (-0.7) 10.3 (+4.0)
 +ACM-8B 15.8 (+0.2) 59.5 (+0.6) 88.0 (+0.2) 9.3 (+3.0)
InternVL2.5-8B 12.8 51.6 84.2 7.1
 +ACM-2B 13.8 (+1.0) 55.7 (+4.1) 84.7 (+0.5) 10.1 (+3.0)
 +ACM-8B 14.8 (+2.0) 57.1 (+5.5) 86.9 (+2.7) 10.2 (+3.1)
Qwen2-VL-7B 13.1 53.5 82.9 6.6
 +ACM-2B 14.6 (+1.5) 55.8 (+2.3) 85.0 (+2.1) 10.6 (+4.0)
 +ACM-8B 15.4 (+2.3) 57.3 (+3.8) 86.7 (+3.8) 10.0 (+3.4)
To validate the generalizability of ACM, we evaluate its performance when integrated with various base models on standard public benchmarks for image, video, and audio captioning. Results show that ACM serves consistently as an effective plug-in, improving controllability and factual alignment.
Image. We report performance on MIA-Bench [22] in Table 8, which focuses on multi-instruction image captioning. Incorporation of ACM consistently improves performance across various backbones, including InternVL2.5 and Yi-VL [64], with particularly notable gains in control-related aspects. With the addition of ACM-8B, we observe significant performance improvements. Moreover, proprietary models such as GPT-4o also benefit from ACM, achieving new state-of-the-art scores, highlighting its plug-and-play potential.
Video. As shown in Table 8, we evaluate video captioning performance using VidCapBench [23], which evaluates not only caption fluency but also key text-to-video attributes such as video aesthetics, content, motion, and adherence to physical laws. Table 8 shows that augmenting InternVL2.5-8B with ACM significantly improves accuracy ($+2.0$), precision ($+5.5$), and conciseness ($+3.1$). These improvements suggest that ACM not only corrects hallucinations but also improves the caption’s capture of crucial video details.
Audio. Controllable audio captioning is less explored. On existing benchmarks [65, 61] (Table 10), despite evaluation metrics being unable to reliably capture semantic-level consistency and the references being relatively short, ACM consistently boosts generation quality, indicating that ACM improves the semantical controllability and fluency of audio captioning.
Table 9: Video generation evaluation. Abbreviations: VQ (Visual Quality), TC (Temporal Consistency), DD (Dynamic Degree), TVA (Text–Video Alignment), and FC (Factual Consistency).
Model VQ TC DD TVA FC Avg
Wan2.1-T2V-14B [66] 2.84 2.78 2.66 2.76 2.77 2.74
 +ACM-8B 3.19 (+0.35) 3.06 (+0.28) 3.18 (+0.52) 3.07 (+0.31) 3.00 (+0.23) 3.10 (+0.36)
HunyuanVideo [67] 2.98 2.89 2.94 2.90 2.83 2.91
 +ACM-8B 3.61 (+0.63) 3.48 (+0.59) 3.62 (+0.68) 3.34 (+0.44) 3.46 (+0.63) 3.50 (+0.59)
Table 10: Audio evaluation results on the AudioCaps and Clotho datasets. Abbreviations: SP(SPICE), SB(Sentence-BERT).
Model AudioCaps Clotho
SP SB SP SB
GPT-4o 0.06 0.41 0.07 0.37
 +ACM-2B 0.06 (+0.00) 0.45 (+0.04) 0.07 (+0.00) 0.40 (+0.03)
 +ACM-8B 0.07 (+0.01) 0.45 (+0.04) 0.08 (+0.01) 0.40 (+0.03)
4.4 Ablation Study
Comparison with SFT, DPO, and self-critic methods. We compare ACM with three widely-used baseline methods on the image modality benchmark: supervised fine-tuning (SFT), direct preference optimization (DPO), and self-critic (SC) approaches. While these baselines require costly re-training for each individual model, ACM provides consistent performance gains without additional model-specific training, demonstrating substantial efficiency advantages. As shown in Table 5, ACM consistently achieves comparable or superior controllability, highlighting its practical benefit of lightweight adaptability.
Ablation on orthogonal data ratios. We conduct orthogonal experiments to explore optimal training data ratios among three data types: ($q$, uncontrolled $a$, $c$), ($q$, $c$, $c$), and ($q$, hallucinated $a$, $c$). Orthogonal designs effectively isolate each factor’s contribution, enabling clear insights with limited experiments. Results in Table 6 suggest that moderate incorporation of slightly incorrect captions ($q,a,c$) improves the model’s sensitivity to subtle instruction deviations. However, determining the globally optimal data ratio remains challenging and demands significant additional training resources. Further details can be found in the Appendix C.
Impact of ($q,c,c$) data proportion. In Fig. 4, we investigate varying proportions of fully compliant ($q,c,c$) data in training. Increasing ($q,c,c$) from $0\%$ to approximately $40\%$ steadily improves performance, yet surpassing $50\%$ leads to a notable decline. This trend implies that moderate exposure to fully correct captions helps the model discern necessary corrections without overfitting. Importantly, ACM thus learns to make nuanced control decisions, reinforcing the robustness and credibility of its caption editing decisions.
4.5 Downstream Utility
We further evaluate ACM’s utility in enhancing noisy captions for downstream video and image generation. Applying ACM to refine captions, including those from the Panda video dataset [68], we observe that the resulting video generations demonstrate improved visual-semantic grounding and alignment compared to those using the original, unrefined captions. Quantitative results for video generation are presented in Table 10; detailed qualitative results for both modalities are in Appendix C. These findings underscore the effectiveness of ACM in real-world multimodal generation pipelines.

### v2

@@ 3.3 AnyCapEval
Figure 3: Details of AnyCapEval. (a) Examples for content evaluation via KPD and style evaluation via scoring rules. (b) AnyCapEval judgments highly align with human preference. (c) Integrating AnyCap consistently boosts base models across both content and style dimensions.
We design AnyCapEval for reliable evaluation of caption quality and, more importantly, the degree of instruction alignment. Beyond previous works, our key idea is that diverse instructions can be grouped into two dimensions: content or style requirements. Accordingly, the evaluation targets: (i) Content, which checks whether captions follow the requested control and remain semantically relevant; and (ii) Style, which measures consistency with human references in narrative manner, structure, and fluency. They are separately evaluated to prevent mutual conflicts, for example when polished language conceals factual mistakes.
Content evaluation. Given a ground-truth caption $c$, a predicted caption $y_{c}$, and a user instruction $q$, we first annotate a keypoint set $\mathcal{K}_{r,q}=\{k_{1},k_{2},\dots,k_{n}\}$ that exhaustively covers the information required by $q$. GPT-4o then serves as an automatic matcher to identify which subset of keypoints is present in $y_{c}$ (see Fig. 3(a) for examples). Different from naive approaches that input $(y_{c},c,q)$ into an LLM to first find keypoints and then produce a score, we constrain the LLM to a binary classification task, i.e., whether $y_{c}$ expresses each $k_{j}$. It simplifies evaluation and improves reliability.
To process verbosity issue, where longer captions may trivially cover more keypoints, we introduce the keypoint density (KPD) to normalize for length, formulated as $\text{KPD}(y_{c};q)=N_{\text{match}}/L_{\text{words}}(y_{c})\times 100$. $N_{\text{match}}$ is the number of matched keypoints and $L_{\text{words}}(y_{c})$ is the word count of $y_{c}$. KPD quantifies the effective information rate, rewarding captions that satisfy user instructions while penalizing redundant or irrelevant content.
@@
Style evaluation. We employ GPT-4o to compare the predicted caption $y_{c}$ with the ground truth $c$ under instruction $q$, producing a discrete score $s(y_{c})\in\{0,\dots,4\}$. The scoring criteria is as follows: (0) severely deviates from $q$ or largely hallucinated/incorrect; (1) deviates from $q$ or contains many hallucinations; (2) slightly deviates from $q$ or contains certain hallucination; (3) highly similar to $c$ and hallucination-free; and (4) outperforms $c$ while remaining hallucination-free; (see Fig. 3(a)).
This scoring follows explicit criteria along three aspects: semantic similarity, hallucination severity, and stylistic conformity, for each level from 0 to 4. With such a constrained rubric, carefully designed prompts, and side-by-side comparison, the evaluation reduces the judge’s degrees of freedom, thereby lowering variance and improving reproducibility.
To verify the bias of AnyCapEval, we conduct a human validation. After assessing multiple models with AnyCapEval, we randomly sample 200 examples per model and ask five independent annotators to judge if the predicted scores and grading rationals (e.g., Fig. 3(a)) aligned with human preference. The results in Fig. 3(b) show high agreement. We also observe that integrating AnyCap consistently improved strong base models in both content and style dimensions (Fig. 3(c)).
@@ 4 Experiments
We train AnyCap in two variants (2B and 8B) on our proposed AnyCapData. Because AnyCap is a relatively small plug-and-play model, it only requires 6 hours on 32 NVIDIA A100 (80GB) GPUs for the 2B variant and 21 hours for the 8B variant to converge. Implementation and training details are provided in Sec. C.1; with a hyperparameter summary in Table 19.
4.1 Evaluation on Instruction-Aligned Captioning
Table 2: Instruction-based image captioning on AnyCapEval. Abbreviations like IPos. are varying instruction types in Table 1. Content and style values are computed using KPD and style scores from Sec. 3.3, respectively. AnyCap significantly improves content accuracy and stylistic fidelity.
Model Content $\uparrow$ Style $\uparrow$
@@ 4.1 Evaluation on Instruction-Aligned Captioning
+AnyCap-2B 3.22 (+1.71) 4.49 (+1.06) 4.82 (+0.28) 4.00 (+1.32) 4.13 (+1.09) 2.27 (+0.14) 1.78 (-0.14) 2.38 (+0.44) 2.71 (+0.63) 2.67 (+0.15) 2.36 (+0.24)
+AnyCap-8B 3.41 (+1.90) 4.82 (+1.39) 4.91 (+0.37) 3.89 (+1.21) 4.26 (+1.22) 2.33 (+0.20) 2.10 (+0.18) 2.56 (+0.62) 2.59 (+0.51) 2.83 (+0.31) 2.46 (+0.34)
Qwen2.5VL-7B 1.51 3.52 5.18 3.03 3.31 2.00 2.02 2.50 2.14 2.63 2.22
+AnyCap-2B 3.38 (+1.87) 4.42 (+0.90) 4.94 (-0.24) 4.68 (+1.65) 4.36 (+1.05) 2.40 (+0.40) 2.04 (+0.02) 2.50 (+0.00) 2.94 (+0.80) 2.72 (+0.09) 2.50 (+0.28)
+AnyCap-8B 3.47 (+1.96) 4.45 (+0.93) 5.84 (+0.66) 4.34 (+1.32) 4.53 (+1.22) 2.36 (+0.36) 2.20 (+0.18) 2.82 (+0.32) 2.88 (+0.74) 2.74 (+0.11) 2.56 (+0.34)
Table 3: Instruction-based video captioning on AnyCapEval. (More details and model comparisons are in Table 27.) AnyCap substantially enhances content adherence and style fidelity.
Model Content $\uparrow$ Style $\uparrow$
@@
+AnyCap-2B 4.67 (+1.59) 6.32 (+1.99) 3.19 (-0.25) 6.49 (+3.33) 4.50 (+1.95) 5.07 (+1.55) 1.84 (+0.45) 1.80 (+0.03) 2.48 (+0.57) 2.39 (+0.05) 2.19 (+0.26)
+AnyCap-8B 4.95 (+1.87) 6.39 (+2.06) 7.03 (+3.59) 5.66 (+2.50) 4.95 (+2.40) 5.73 (+2.21) 2.00 (+0.61) 1.88 (+0.11) 2.20 (+0.29) 2.61 (+0.27) 2.24 (+0.31)
Qwen2.5VL-7B 2.88 4.20 3.29 2.24 2.93 3.54 1.63 1.65 2.40 2.29 2.10
+AnyCap-2B 3.39 (+0.51) 5.97 (+1.77) 6.23 (+2.94) 6.30 (+4.06) 4.48 (+1.55) 5.25 (+1.71) 1.66 (+0.03) 1.82 (+0.17) 2.30 (-0.10) 2.55 (+0.26) 2.16 (+0.06)
+AnyCap-8B 4.03 (+1.15) 6.15 (+1.95) 6.58 (+3.29) 6.04 (+3.80) 4.82 (+1.89) 5.55 (+2.01) 1.92 (+0.29) 1.98 (+0.33) 2.45 (+0.05) 2.55 (+0.26) 2.31 (+0.21)
Table 4: Instruction-based audio captioning on AnyCapEval. AnyCap consistently improves both content and style metrics over the corresponding base captioners.
Model Content $\uparrow$ Style $\uparrow$
@@
+AnyCap-2B 1.55 (+0.13) 1.37 (+0.35) 1.32 (+0.64) 0.92 (+0.49) 1.21 (+0.50)
+AnyCap-8B 1.48 (+0.06) 1.23 (+0.21) 1.29 (+0.61) 1.13 (+0.70) 1.22 (+0.51)
Table 5: Performance on MIA-Bench. Metrics include accuracy (desc), mention (ment), length (len), perspective (persp), genre (gen).
Model desc. $\uparrow$ ment. $\uparrow$ len. $\uparrow$ persp. $\uparrow$ gen. $\uparrow$ avg. $\uparrow$
GPT-4o 90.3 87.9 90.5 85.2 91.6 89.1
 +AnyCap-2B 89.1 (-1.2) 88.9 (+1.0) 90.3 (-0.2) 90.0 (+4.8) 92.2 (+0.6) 90.1 (+1.0)
 +AnyCap-8B 90.4 (+0.1) 89.4 (+1.5) 89.9 (-0.6) 89.2 (+4.0) 92.9 (+1.3) 90.3 (+1.2)
InternVL2.5-8B 85.2 79.0 82.6 72.5 85.2 80.9
 +AnyCap-2B 85.7 (+0.5) 79.6 (+0.6) 85.2 (+2.6) 77.5 (+5.0) 85.7 (+0.5) 82.7 (+1.8)
 +AnyCap-8B 86.9 (+1.7) 79.0 (+0.0) 86.1 (+3.5) 88.9 (+16.4) 86.3 (+1.1) 85.4 (+4.5)
Yi-VL-34B 71.7 52.0 55.8 60.0 59.8 59.9
 +AnyCap-2B 71.8 (+0.1) 54.2 (+2.2) 61.9 (+6.1) 61.7 (+1.7) 61.0 (+1.2) 62.1 (+2.2)
 +AnyCap-8B 79.5 (+7.8) 66.5 (+14.5) 67.1 (+11.3) 63.0 (+3.0) 75.8 (+16.0) 70.4 (+10.5)
Table 6: Performance on AudioCaps and Clotho. Metrics: spice (sp), sentence-bert (sbt).
Model AudioCaps $\uparrow$ Clotho $\uparrow$
sp. sbt. sp. sbt.
GPT-4o 0.06 0.41 0.07 0.37
+AnyCap-2B 0.06 (+0.00) 0.45 (+0.04) 0.07 (+0.00) 0.40 (+0.03)
+AnyCap-8B 0.07 (+0.01) 0.45 (+0.04) 0.08 (+0.01) 0.40 (+0.03)
Table 7: Performance on VidCapBench. Metrics include accuracy (acc), precision (pre), coverage (cov), and conciseness (con).
Model acc. $\uparrow$ pre. $\uparrow$ cov. $\uparrow$ con. $\uparrow$
GPT-4o 15.6 58.9 87.8 6.3
+AnyCap-2B 15.4 (-0.2) 59.1 (+0.2) 87.1 (-0.7) 10.3 (+4.0)
+AnyCap-8B 15.8 (+0.2) 59.5 (+0.6) 88.0 (+0.2) 9.3 (+3.0)
InternVL2.5-8B 12.8 51.6 84.2 7.1
+AnyCap-2B 13.8 (+1.0) 55.7 (+4.1) 84.7 (+0.5) 10.1 (+3.0)
+AnyCap-8B 14.8 (+2.0) 57.1 (+5.5) 86.9 (+2.7) 10.2 (+3.1)
Qwen2VL-7B 13.1 53.5 82.9 6.6
+AnyCap-2B 14.6 (+1.5) 55.8 (+2.3) 85.0 (+2.1) 10.6 (+4.0)
+AnyCap-8B 15.4 (+2.3) 57.3 (+3.8) 86.7 (+3.8) 10.0 (+3.4)
@@
We evaluate the impact of integrating AnyCap with various powerful base models (proprietary models, e.g., GPT-4o (Hurst et al., 2024); open-source models, e.g., InternVL2.5 (Chen et al., 2024c), Qwen2.5-VL (Bai et al., 2025), MiniCPM-o (Yao et al., 2024)) using our AnyCapEval benchmark across the image, video, and audio modalities. As shown in Tables 2, 3 and 7, adding AnyCap results in significant improvements in instruction alignment across all modalities and base models, particularly enhancing content fidelity. Fig. 3(c) visualizes these gains for GPT-4o and InternVL2.5-8B.
@@
Model-centric analysis. The results consistently show the effectiveness of AnyCap in improving instruction alignment ability. Both the 2B and 8B variants significantly enhance instruction adherence compared to the unassisted base models across all modalities, with the larger AnyCap-8B generally yields greater improvements. Notably, with AnyCap-8B, certain open-source models achieve instruction alignment scores that are comparable or even exceed strong unassisted proprietary baselines. For example, AnyCap-8B elevates the performance of open-source InternVL2.5 from $\sim 2.1$ to $\sim 2.5$ in the image style category. The enhanced score surpasses the $\sim 2.3$ achieved by GPT-4o.
@@
Task-centric analysis. Boosts in content-related instruction alignment typically surpass those in style, likely because content controls provide more explicit learning signals, whereas stylistic constraints are inherently more subjective and harder to enforce. Visual tasks outperform audio counterparts, potentially due to the richer supervision and higher information density in vision-language pretraining. AnyCap exhibits varied effects across settings: in well-studied tasks (e.g., image appearance), it refines already strong captions; in less-explored, complex tasks (e.g., video actions or audio events), it fills in missing content, yielding larger gains.
@@ 4.2 Evaluation on Public Benchmarks
To assess AnyCap’s generalizability, we integrate it with diverse backbones and evaluate on public image, video, and audio captioning benchmarks. On MIA-Bench (Qian et al., 2025) (Table 7), AnyCap consistently improves performance across models such as InternVL2.5 and Yi-VL (Young et al., 2024). When scaled to AnyCap-8B, it achieves further gains, and even proprietary models such as GPT-4o benefit, reaching new state-of-the-art results. On VidCapBench (Chen et al., 2025b) (Table 7), augmenting InternVL2.5-8B with AnyCap yields improvements in accuracy (+2.0), precision (+5.5), and conciseness (+3.1), indicating fewer hallucinations and better capture of motion and content cues. For audio, although references are shorter and metrics more sensitive on Clotho and AudioCaps (Drossos et al., 2020; Kim et al., 2019) (Table 7), AnyCap still consistently enhances fluency and semantic controllability. Collectively, AnyCap reliably steers strong backbones toward more factually grounded captions across modalities.
4.3 Component-wise Analysis
Table 8: AnyCap vs. SFT, DPO, SC on AnyCapEval. Entries are relative improvements (%) over each model’s origin baseline under different paradigms, reported separately for content (Cont) and style (Sty). AnyCap demonstrates clear effectiveness. Model (Modality) SFT DPO SC AnyCap-2B AnyCap-8B Cont. Sty. Cont. Sty. Cont. Sty. Cont. Sty. Cont. Sty. InternVL2.5-8B(Ima) +9.2 +3.5 +9.5 +4.4 +3.3 +0.5 +35.9 +11.3 +40.1 +16.0 InternVL2.5-8B(Vid) +2.3 +1.0 +2.8 +1.6 +0.0 -1.6 +44.0 +13.5 +62.8 +16.1 Qwen2.5VL-7B(Ima) +6.7 +3.8 +7.5 +4.1 +11.5 +5.0 +31.7 +12.6 +36.9 +15.3 Qwen2.5VL-7B(Vid) +3.2 +3.0 +7.3 +5.2 +16.7 +11.0 +26.6 +2.9 +56.8 +10.0 Table 9: AnyCap vs. SFT, DPO, SC on MIA benchmark with InternVL2.5-2B. SFT DPO SC AnyCap 1.2% +7.0% +2.8% +16.9% Table 10: Data ratio comparison with InternVL2.5-8B. 1:1:1 1:2:2 2:1:2 2:2:1 +17.9% +17.2% +16.5% +21.1%
Figure 4: Impact of training data ratio.
Figure 5: Human evaluation comparing AnyCap-8B with GPT-4o. Captions refined with AnyCap align more closely with the given instructions, enhancing both content and style consistency.
Comparison with SFT, DPO, and Self-Critic approaches. One motivation for plug-and-play AnyCap is to avoid retraining each base model. We compare it against three options: (i) supervised fine-tuning (SFT), where instructions are added to prompts to fine-tune the base model; (ii) direct preference optimization (DPO) on base models using preferred–non-preferred pairs from AnyCapData; and (iii) a self-critic (SC) approach where the backbone revises its own outputs at test time. Across both AnyCapEval (Sec. 4.3) and MIA (Table 9), AnyCap consistently achieves stronger performance. This advantage arises as it builds on base captions as references, focusing learning on refining outputs toward user instructions rather than generating fully aligned captions from scratch.
Ablation on training data ratio. Our triplet training data consist of three main types: ($q$, uncontrolled $a$, $c$), ($q$, $c$, $c$), and ($q$, hallucinated $a$, $c$). To explore the optimal ratio, we first vary the share of fully correct samples ($q$, $c$, $c$). As shown in Fig. 5, the average accuracy on three modalities improves steadily up to around $40\%$, but declines when exceeding $50\%$. This suggests that although exposure to correct captions benefits learning, excessive reliance may lead to overfitting.
We further perform ablation across all three types. Results in Table 10 show that including more uncontrolled $a$ samples yields better performance than emphasizing hallucinated $a$. A likely reason is that outputs from strong base models are more prone to instruction violations than factual errors, so AnyCap gains more from frequently correcting uncontrolled cases.
Keypoint coverage v.s. caption length. In Table 11, we analyze the relationship between caption length and keypoint coverage. An important property of AnyCap is that it does not “win by writing more”. Instead, it effectively reduces unrelated details from base models (especially GPT-4o, as visualized in Fig. 6) while maintaining or even improving keypoint accuracy. The ability to generate concise yet accurate captions is one reason why AnyCap achieve higher scores on AnyCapEval.
Table 11: Absolute keypoint counts (no length normalization) and response lengths on AnyCapEval image captioning. Model Keypoints (Abs.) $\uparrow$ Response Length Per. Ins. IApp. IPos. Per. Ins. IApp. IPos. GPT-4o 2.96 3.52 3.13 1.42 95.8 91.8 92.4 93.7 +AnyCap 2.29 4.00 3.13 1.55 49.6 83.3 61.7 51.8 InternVL2.5-8B 2.54 3.64 2.85 1.21 90.0 84.6 83.0 68.0 +AnyCap 2.44 3.76 2.85 1.70 62.5 76.6 59.4 49.8 Qwen2.5VL-7B 2.68 3.80 2.82 1.21 84.0 70.3 79.3 74.2 +AnyCap 2.59 3.78 2.95 1.64 64.0 68.0 66.7 51.6 Table 12: Human evaluation on AnyCapEval comparing manually written captions and AnyCap. AnyCap matches non-experts but remains slightly below expert performance, indicating clear headroom for future improvement. Image $\uparrow$ Video $\uparrow$ Audio $\uparrow$ Cont. Sty. Cont. Sty. Cont. Sty. Expert 4.62 2.77 6.04 2.62 3.30 4.13 Non-expert 4.42 2.66 5.72 2.35 2.01 1.37 GPT+AnyCap 4.54 2.65 5.74 2.32 1.88 1.30
4.4 Human Evaluation
To complement model-based evaluation, we conduct two-fold human study. First, a preference test (Fig. 5) compare AnyCap-8B with GPT-4o on both visual and auditory modalities, asking annotators which better satisfied content and style criteria. Second, we hire PhD-level humanities experts and well-educated non-experts to handwrite captions for comparison with AnyCap. As shown in Table 12, experts surpassed AnyCap in fine control, while AnyCap matched or slightly outperformed the non-expert group on visual tasks and remained broadly comparable on audio. This suggests that AnyCap delivers strong quality while aligning well with human preferences.